# Optimistic Exploration for Risk-Averse Constrained Reinforcement Learning

### v1

@@ 5.3 Cost-Reward Trade-off Evaluation
Reward Mean Cost $\text{CVaR}_{0.05}$ Unserved Energy in Outage (%) Outage Temp. Violations (%) SAC Lag -605.9$\pm$8.3 772.6$\pm$15.2 861.2$\pm$27.0 0.61$\pm$0.01 0.73$\pm$0.01 WCSAC-0.05 -1417.6$\pm$118.9 623.3$\pm$39.1 699.7$\pm$38.7 0.66$\pm$0.03 0.70$\pm$0.04 ORAC-0.05 -1299.3$\pm$57.4 629.6$\pm$17.6 702.8$\pm$25.9 0.63$\pm$0.02 0.65$\pm$0.03
To assess whether or not ORAC improves the reward-cost trade-off over the baseline risk-averse WCSAC by converging to higher reward policies, we evaluate the algorithms using Safety-Gymnasium environments PointGoal1 and and PointButton1 [18], and with complex building energy management simulation environment, CityLearn [30].
@@
In our experiments we represent the goal of reducing the building’s demand on the energy grid through a reward function and the goal of maintaining the indoor temperature through a cost function and constraint limit, representing a comfort band around the desired temperature. We set this comfort band constraint to 1 degree at every timestep. Each episode runs for a total of 720 timesteps, representing 1 month. We fix the probability of a power outage at approximately 10% at each timestep, which may appear high but outages occur for consecutive timesteps and their impact may vary depending on a number of factors. We train our agents for 3 million timesteps at a risk-level of 0.05, across 10 seeds. We measure the mean episodic reward and cost, and the $\text{CVaR}_{0.05}$ episodic costs. We include two metrics specific to CityLearn, Outage Temperature Violations indicating the proportion of timesteps that the agent violates the temperature constraint specifically during an outage, and Unserved Energy during Outages indicating the proportion of energy demand that is unmet by the agent specifically during an outage. Both these metrics highlight the agents performance specifically during outages as the reward and cost metrics are calculated during both regular operation and during outages.
Table 3 shows the results of our evaluations. As expected, the risk-neutral basline, SAC Lag, achieves the highest episodic reward, but returns the highest mean and CVaR cost. The risk-averse baseline, WCSAC, achieves the lowest cost returns but the lowest mean episodic reward. ORAC improves significantly on the mean episodic reward over WCSAC, but returns slightly higher mean and CVaR costs. We note however, that ORAC significantly improves upon both the two baselines in terms of Outage Temperature Violations and achieves an Unserved Energy during Outages score close to SACLag, which indicates better exploitation of reward during outages that WCSAC. We reason that these results along with the increased cost shown by ORAC may indicate a qualitative change in the policy by focusing more on maintaining temperature demands and energy supply during outages, rather than during regular hours.
@@ 6 Discussion
In this work we presented ORAC, an optimistic exploration approach to Risk-averse Constrained RL for Lagrangian based off-policy actor-critic algorithms. Specifically, ORAC constructs an optimistic exploratory policy at each timestep, that maximises an upper confidence bound of the state-action reward value function whilst minimising a local lower confidence bound of the risk-averse state-action cost value function. ORAC is encouraged to explore uncertain regions of the environment, to discover high reward policies whilst aiming to still satisfy safety constraints. We demonstrated that ORAC, through improved exploration, prevents a risk-averse baseline algorithm, WCSAC, from converging to sub-optimal policies that fail to find the correct solution in a risky GridWorld environment. We further demonstrated improvements in Safety-Gymnasium environments and in a complex building energy management environment, CityLearn, ORAC improved reward maximisation whilst minimising cost, over baseline algorithms such as WCSAC.
Similarly to the underlying algorithm, WCSAC, however, ORAC does not guarantee safe exploration. In future work we aim to investigate an exploration approach that adaptively explores the environment switching between optimism and conservatism based on constraint satisfaction, to improve the safety during training. Further future could look at most robust estimates of epistemic uncertainty and potentially dynamic confidence bounds, as ORAC makes use of fixed hyperparameter values for the confidence bounds.
@@ B.2 Spectral Risk - Linear Approximation
[…]
@@
[…]
@@ B.2.1 Optimistic Lower Bound
[…]
@@ Appendix C Comparison of IQN vs Spectral Risk - Linear Approximation
PointGoal1 PointButton1 Reward Mean Cost $\text{CVaR}_{0.5}$ $\text{CVaR}_{0.25}$ Reward Mean Cost $\text{CVaR}_{0.5}$ $\text{CVaR}_{0.25}$ WCSAC-IQN 34.9$\pm$0.5 8.0$\pm$2.6 14.3$\pm$4.2 19.7$\pm$6.4 16.2$\pm$4.9 14.8$\pm$3.4 25.8$\pm$5.7 35.9$\pm$7.0 WCSAC-Lin 33.9$\pm$0.5 5.1$\pm$1.0 9.0$\pm$1.6 12.3$\pm$1.9 14.6$\pm$3.5 7.2$\pm$2.5 13.0$\pm$4.4 17.6$\pm$5.3

### v2

@@ 5.3 Cost-Reward Trade-off Evaluation
[…]
To assess whether or not ORAC improves the reward-cost trade-off over the baseline risk-averse WCSAC by converging to higher reward policies, we evaluate the algorithms in the Safety-Gymnasium environments PointGoal1 and PointButton1 [18], and with complex building energy management simulation environment, CityLearn [31].
@@
In our experiments we represent the goal of reducing the building’s demand on the energy grid through a reward function and the goal of maintaining the indoor temperature through a cost function and constraint limit, representing a comfort band around the desired temperature, which is set to 1 degree at every timestep. Each episode runs for a total of 720 timesteps, representing 1 month. We fix the probability of a power outage at approximately 10% at each timestep, which may appear high but outages occur for consecutive timesteps meaning their impact varies, depending on a number of factors. We train our agents for 3 million timesteps at a risk-level of 0.05, across 10 seeds. We measure the mean episodic reward and cost, and the $\text{CVaR}_{0.05}$ episodic costs. We include two metrics specific to CityLearn, Outage Temperature Violations indicating the proportion of timesteps that the agent violates the temperature constraint during an outage, and Unserved Energy during Outages indicating the proportion of energy demand that is unmet by the agent specifically during an outage. These metrics highlight the agents performance specifically during outages as the reward and cost metrics are calculated during both regular operation and during outages.
Table 3 shows the results of our evaluations. As expected, the risk-neutral basline, SAC Lag, achieves the highest episodic reward, but returns the highest mean and CVaR cost. The risk-averse baseline, WCSAC, achieves the lowest cost returns but the lowest mean episodic reward. ORAC improves significantly on the mean episodic reward over WCSAC, but returns slightly higher mean and CVaR costs, still satisfying the threshold. ORAC significantly improves upon both baselines in terms of Outage Temperature Violations and achieves an Unserved Energy during Outages score close to SACLag, which indicates better exploitation of reward during outages than WCSAC. We reason that these results along with the increased cost shown by ORAC may indicate a qualitative change in the policy by focusing more on maintaining temperature demands and energy supply during outages, rather than during regular hours.
@@ 6 Discussion
In this work we presented ORAC, an optimistic exploration approach to Risk-averse CRL for Lagrangian based off-policy actor-critic algorithms. Specifically, ORAC constructs an optimistic exploratory policy at each timestep, that maximises an upper confidence bound of the state-action reward value function whilst minimising a local lower confidence bound of the risk-averse state-action cost value function. ORAC is encouraged to explore uncertain regions of the environment, to discover high reward policies whilst aiming to still satisfy safety constraints. We demonstrated ORAC, by improved exploration, prevents a risk-averse baseline, WCSAC, from converging to sub-optimal policies that fail to find the correct solution in a risky GridWorld environment. We further demonstrated improvements in Safety-Gymnasium environments and in a complex building energy management environment, CityLearn, where ORAC improved reward maximisation whilst minimising cost, over the baseline.
Similarly to the underlying algorithm, WCSAC, however, ORAC does not guarantee safe exploration but instead aims for safe policies at deployment. In future work we aim to investigate an exploration approach that adaptively explores the environment switching between optimism and conservatism based on constraint satisfaction, to improve the safety during training. Further future work could look at most robust estimates of epistemic uncertainty and potentially dynamic confidence bounds, as ORAC makes use of fixed hyperparameter values for the confidence bounds.
[…]
The author would like to acknowledge the support of the IBM Pre-Doc Programme, part-funded by IDA Training Grant 215544.
@@ B.2 Spectral Risk - Linear Approximation
$\mathcal{R}_{\sigma}(X)\vcentcolon=\int_{0}^{1}F_{X}^{-1}(u)\sigma(u)du$ (29)
where $\sigma$ is an increasing function $\sigma\geq 0,\int_{0}^{1}\sigma(u)du=1$. The dual form of E.q. 29 can be expressed by:
$\mathcal{R}_{\sigma}(X)=\inf_{g}\mathrm{E}[g(X)]+\int^{1}_{0}g^{*}(\sigma(u))du$ (30)
@@
$\mathcal{R}_{\hat{\sigma}}^{\beta}\vcentcolon=\mathbb{E}[g_{\beta}(X)]+\int_{0}^{1}g^{*}(\hat{\sigma}(u))du$ (34)
@@ B.2.1 Optimistic Lower Bound
$\hat{R}_{\sigma}^{\beta}=\frac{1}{K}\sum_{k=0}^{K}\bigg{[}\eta_{1}\hat{q}_{C}^{k}(s,a)+(\eta_{2}-\eta_{1})(\hat{q}_{C}^{k}(s,a)-\beta)_{+}\bigg{]}+\int_{0}^{1}g^{*}(\hat{\sigma}(u))du$ (38)
@@ Appendix C Comparison of IQN vs Spectral Risk - Linear Approximation
[…]
Appendix D Ablations
ORAC introduces a small number hyperparameters that each affect its performance. To assess the impact of each value we train ORAC under various hyperparameter values, namely the $\beta$ values used to calculate the upper and lower bounds on the reward and cost function, respectively, and the size of the cost function ensemble. We train ORAC in the PointButton environment at CVaR 0.5, following the main experiments, varying the value for each of these hyperparameters and assess their respective impact on the performance on ORAC.
D.1 Optimistic Lower Bound - Cost Function
We begin by establishing the impact of varying the value for the optimistic lower bound, $\beta_{C}$, of the state-action cost quantile function. Recall that the locally optimistic lower bound state-action cost value is calculated through Eq. 15:
$\hat{q}_{C}^{k}(s,a)=\mu_{q_{C}}^{k}(s,a)-\beta_{C}[\sigma_{q_{C}}^{k}(s,a)]$ (39)
Table 7 outlines ablations across 3 values of $\beta_{C}$, results for value 0.5 and 1.5 are reported to 4 seeds whilst results for the value 1.0 are obtained from the main set of experiments. We note that the value used in the main set of experiments, 1, finds the best constraint satisfying solution. Interestingly the lowest value of $\beta_{C}$, 0.5, higher reward than the other values but also achieved higher cost. This might indicate that there is a threshold value, above which optimistic exploration improves the policy’s cost performance by gathering more useful experience that the policy is uncertain about.
Table 7: Results in SafetyGymnasium’s PointButton1, across different values for $\beta_{C}$. Safety threshold set to 10 for all environments.
[…]
D.2 Optimistic Upper Bound - Reward Function
Next we evaluate the impact of the optimistic uppper bound value $\beta_{R}$ of the state-action reward function. Recall that the locally optimistic upper bound state-action reward value is calculated through Eq. 14:
$\hat{Q}_{R}^{\pi}(s,a)=\mu_{Q_{R}}(s,a)+\beta_{R}[\sigma_{Q_{R}}(s,a)]$ (40)
Table 8 outlines ablations across 3 values of $\beta_{R}$, results for value 2 and 5 are reported to 4 seeds whilst results for the value 4 are obtained from the main set of experiments. We note again the value, 4, used in the main set of experiments finds the best solution, satisfying the constraint whilst only slightly reducing reward compared to WCSAC. The value 2 shows lower reward than the value 4, indicating ORAC is not optimistic enough in terms of reward. The value 5 improves slightly in terms of reward over the value 4, but achieves higher cost. This trend indicates similarly to the lower bound value $\beta_{C}$, that there is a sweet spot value that maximises reward whilst satisfying the cost constraint.
Table 8: Results in SafetyGymnasium’s PointButton1, across different values of $\beta_{R}$. Safety threshold set to 10 for all environments.
[…]
D.3 Cost Function Ensemble Size
To evaluate the performance impact of increasing the number of cost functions that make up the ensemble used to train both the baseline WCSAC and ORAC, we vary the size of the ensemble and detail the changes in episodic reward and cost, and CVaR costs. We trained 4 seeds of each algorithm is trained for 5 million timesteps.
ORAC outperforms WCSAC in terms of reward with an ensemble larger than 2 cost functions in the more challenging CVaR 0.25 experiments. When paired with an ensemble of 8 cost functions, ORAC performs better than the main experimental setup of in the challenging CVaR 0.25, but worse in the CVaR 0.5 setting. Otherwise, WCSAC and ORAC with 5 ensembles performs the best consistently across the two experiment settings.
Table 9: Results in SafetyGymnasium’s PointButton1, across different cost function ensemble sizes. Safety threshold set to 10 for all environments.
[…]